A -17nC charge is located at the origin.

1. What is the strength of the electric field at the position (x,y)=(0cm,5.0cm)?

Express your answer using two significant figures.

2. What is the strength of the electric field at the position (x,y)=(?5.0cm,?5.0cm)?

Express your answer using two significant figures

3. What is the strength of the electric field at the position (x,y)=(?5.0cm,5.0cm)?

Express your answer using two significant figures
Answer #1

A) E = KQ/R^2

= -17*10^-9 * 9*10^9 / 0.05/0.05

= -61200 i

b) E = -61200i -61200j

c) E = -61200i +61200j
Answer #2

we have \bar{E}=\frac{Kq}{r^{2}}\hat{r} where K=9*10^{9}Nm^{2}C^{-2} and q=-17nC

hence

1.\bar{E}=\frac{Kq}{r^{2}}\hat{r}=6.12*10^{4}V/m(-\hat{i})

2.\bar{E}=\frac{Kq}{r^{2}}\hat{r}=2.16*10^{4}V/m(\hat{i}+\hat{j}) here r=5\sqrt{}2

3.\bar{E}=\frac{Kq}{r^{2}}\hat{r}=2.16*10^{4}V/m(\hat{i}-\hat{j})here also r=5\sqrt{}2